A violinist places her finger so that the vibrating section of a 1.00 g/m string has a length of 40.0 cm, then she draws her bow across it. A listener nearby in a 20 degrees Celcius room hears a note with a wavelength of 60.0 cm. What is the tension in the string?
Answer #1
first find the frequency heard...
speed of sound at 20 deg C = 343 m/s
   f = v / λ = 343 /0.60   = 571.7 Hz
Then... find the wavelength of the wave on the string:
   λ = 2L = 2 * 0.40 = 0.80 meters
Find the speed of the waves on the string:   speed = λ f = 0.80 * 571.7 =    457.3 m/s
Then use:     speed2 = Tension / μ     or
   Tension = speed2 *μ = 457.32 * 0.001 =    209.2 Newtons